OpenStudy (anonymous):

What value of x makes the equation true? –7.12 = –4.8 + x

OpenStudy (anonymous):

@iGreen

sammixboo (sammixboo):

Well to solve this all you need to do is subtract -4.8 on both sides

OpenStudy (igreen):

Add 4.8 to both sides, what's -7.12 + 4.8?

OpenStudy (anonymous):

Is it add or subtract

OpenStudy (igreen):

Subtracting -4.8 is the same thing as adding 4.8.

sammixboo (sammixboo):

of the equation -7.12 - (-4.8) = -4.8 - (-4.8)+ x -7.12 + 4.8= -4.8 + 4.8 + x -7.12 + 4.8 = x

OpenStudy (igreen):

Subtracting(negative) -4.8(negative) Negative and Negative is a positive.

OpenStudy (anonymous):

ok

OpenStudy (igreen):

So just add -7.12 + 4.8.

OpenStudy (anonymous):

-2.3

OpenStudy (anonymous):

yeah i think so

sammixboo (sammixboo):

I will show you a trick 

OpenStudy (igreen):

Yep, -2.3 is your answer.

OpenStudy (anonymous):

thanks everyone

sammixboo (sammixboo):